# Open, Monetizable, and Loyal AI

### v1

@@ Abstract
Artificial Intelligence (AI) has steadily improved across a wide range of tasks, and a significant breakthrough towards general intelligence was achieved with the rise of generative deep models, which have garnered worldwide attention. However, the development and deployment of AI are almost entirely controlled by a few powerful organizations and individuals who are racing to create Artificial General Intelligence (AGI). These centralized entities make decisions with little public oversight, shaping the future of humanity, often with unforeseen consequences.
In this paper, we propose OML, which stands for Open, Monetizable, and Loyal AI, an approach designed to democratize AI development and shift control away from these monopolistic actors. OML is realized through an interdisciplinary framework spanning AI, blockchain, and cryptography. We present several ideas for constructing OML systems using technologies such as Trusted Execution Environments (TEE), traditional cryptographic primitives like fully homomorphic encryption and functional encryption, obfuscation, and AI-native solutions rooted in the sample complexity and intrinsic hardness of AI tasks.
A key innovation of our work is the introduction of a new scientific field: AI-native cryptography, which leverages cryptographic primitives tailored to AI applications. Unlike conventional cryptography, which focuses on discrete data and binary security guarantees, AI-native cryptography exploits the continuous nature of AI data representations and their low-dimensional manifolds, focusing on improving approximate performance. One core idea is to transform AI attack methods, such as data poisoning, into security tools. This novel approach serves as a foundation for OML 1.0, an implemented system that demonstrates the practical viability of AI-native cryptographic techniques. At the heart of OML 1.0 is the concept of model fingerprinting, a novel AI-native cryptographic primitive that helps protect the integrity and ownership of AI models.
The spirit of OML is to establish a decentralized, open, and transparent platform for AI development, enabling the community to contribute, monetize, and take ownership of AI models. By decentralizing control and ensuring transparency through blockchain technology, OML prevents the concentration of power and provides accountability in AI development that has not been possible before.
To the best of our knowledge, this paper is the first to:
  • Identify the monopolization and lack of transparency challenges in AI deployment today and formulate the challenge as OML (Open, Monetizable, Loyal).
  • Provide an interdisciplinary approach to solving the OML challenge, incorporating ideas from AI, blockchain, and cryptography.
  • Introduce and formally define the new scientific field of AI-native cryptography.
  • Develop novel AI-native cryptographic primitives and implement them in OML 1.0, analyzing their security and effectiveness.
  • Leverage blockchain technology to host OML solutions, ensuring transparency, decentralization, and alignment with the goals of democratized AI development.
Through OML, we aim to provide a decentralized framework for AI development that prioritizes open collaboration, ownership rights, and transparency, ultimately fostering a more inclusive AI ecosystem.
Introduction
@@ 1.1 Era of AI
Artificial Intelligence (AI) has steadily improved in a wide range of tasks. Robots now handle household chores, like vacuuming [1, 2]; AI systems outperform humans in games like Chess and Go [3, 4, 5] and in formal mathematical reasoning, such as with Alpha Proof by Google DeepMind [6]. AI has also made significant contributions to scientific research, notably in protein structure prediction [7, 8], advancing drug discovery [9, 10, 11] and, more recently, in the FunSearch program by Google DeepMind [12]. One of the most significant breakthroughs towards general intelligence was the rise of generative deep models, such as GPT4 [13, 14], which garnered worldwide attention. These large language models (LLMs) demonstrate extraordinary proficiency in natural language processing and excel in diverse fields like medicine, law, accounting, computer programming, and music. Moreover, they can efficiently interface with external tools such as search engines, calculators and APIs to perform tasks with minimal guidance, demonstrating their impressive adaptability and learning capability. This breakthrough indicates that AI is on track to dominate and, in many cases, replace human interactions, becoming a crucial innovation engine for all human activities and the way societies organize and govern themselves.
The development and deployment of AI are almost entirely controlled by a few powerful organizations, led by a handful of individuals, who are feverishly racing to create Artificial General Intelligence (AGI). Their decisions – made with little public oversight – will shape the future of humanity, often with unforeseen consequences. At the same time, millions of people are acquiring AI-related skills, striving to contribute to the field. Yet, the concentration of power in AI development leaves them with limited opportunities to showcase their talents, and even fewer chances for obtaining meaningful employment.
Figure 1.1: A representation of the current AI paradigm. On the left, the millions of aspiring AI developers in the world. In the middle, how they are noisily filtered through to compete for employment. On the right, a select few corporations controlling the world’s AI demand and supply.
At Sentient, our mission is to democratize AI development by bringing ownership rights to an open, collaborative platform and enabling community-built AI. We aim to create the science and technology that empowers anyone to build, collaborate, own, and monetize AI products – ushering in a new era of AI entrepreneurship and a community-built open AGI.
Our vision is to establish a vibrant ecosystem where researchers, developers, and users are natively incentivized to collaborate on an open AGI platform, transcending the closed, monolithic AI systems that dominate today. By unlocking the potential of millions of AI developers to contribute to AGI in an open and transparent way, we can ensure that AI development is aligned with the needs of humanity. With more individuals working on building AI that benefits everyone, we gain a broader and more diverse pool of talent. This way, many minds working together can better guide AI toward a future that serves all of humanity.
As a community-built open AGI platform, Sentient will enable a new model of governance – one where the community collectively shapes the development, use, and safety of AGI. This decentralized approach offers a level of transparency and accountability that has never before been possible in AI development, instead of ceding control to a handful of individuals or corporations. We believe the future of the AI economy is open – an ecosystem that fosters both competition and collaboration and fairly rewards contributions. New cryptographic primitives and protocols need to be invented towards this goal – formulating the appropriate scientific agenda and proposing solutions are the primary goals of this manuscript.
1.2 Community-built AI
With true community-built AI, the participants who contributed to the AI have the technological freedom to be rewarded for their contributions and to decide on how it can be used. The AI builders can introduce new models into the ecosystem, fine-tune existing models, provide data or filter data for training models. AI users will download and use models for creating many new AI applications. In fact, even AI users can contribute to AI by providing the data associated with the usage; in this role, they serve as AI builders.
To align the incentives of builders with the growth of the AI economy through innovations, we need to make sure that as more users download and use AI models, the contributors involved are rewarded. In Figure 1.2 we can see how AI platforms work now and the focus of this work, which is to close the loop properly by incentivizing the contributors in AI.
@@ 1.2 Community-built AI
Figure 1.2: On the left, a one-way interaction is depicted, where AI platforms deliver outputs to users without feedback. On the right, the process evolves into a closed-loop system, where user feedback influences the AI platform, aligning incentives and enabling continuous improvement in an open environment.
In particular, we need to make sure that the following requirements are met:
@@
    Open Models. All models are available to be downloaded and used for different AI applications;
@@
    Decentralized Control. The access to a model for downloading and permission for using is decentralized (not controlled by any one party);
@@
    Trust-free Monetization. The assertion of contributions by AI builders and the assignment of corresponding rewards on usage should happen securely, without any additional trust assumptions.
The AI that we see all around us now was built by combining multiple open-source contributions. Until recently, all core libraries were open and powerful models such as BERT [15], and even GPT [13, 14], were available openly. However, as AI matured and the economic potential of this powerful technology became clear, most large companies developing AI in the open have switched to a closed development strategy. These companies geared their efforts towards dominating the AI economy where everyone else will be a high-level user of the AI that they will build [16, 17, 18]. This is not only unfair to the original contributors, who are denied their fair share in this new economy, but hinders AI innovation. Furthermore, this poses challenges of AI safety [19, 20] and alignment [21, 22], which is no longer decided based on broader goals and consensus but is determined by the business and political priorities of the closed AI companies.
In the other direction, open models such as Llama [23] continue to improve and remain at par with closed models. Every industry in the world is adopting AI and the variety of use cases increases by the day. The pool of AI innovators is expanding rapidly and a large part of global work force wishes to contribute to AI. Combining these forces can lead to a much more powerful and inclusive AI, which will be community-built and open for innovations. But this requires a new socio-technical framework for aligning incentives for AI builders, allowing them to fairly share rewards of the AI economy they help build. Our goal in building the Sentient Protocol is to address this problem.
1.3 AI Service Landscape
Today, AI is being delivered to users via two different service models.
  • Closed. In this paradigm, the primary method for accessing AI models is through public inference APIs [16, 17, 18]. For instance, the OpenAI API enables users to interact with models like ChatGPT [24] and DALL-E [25] via a web interface. Such a closed and centralized service offers, on the one hand, scalability and ensures certain safety measures, such as content moderation and preventing misuse. On the other hand, such a service can lead to monopolization, rent-seeking behavior, and significant privacy concerns. Additionally, users have no control over the service they pay for, as the model owners can arbitrarily filter user inputs, alter outputs, or change the underlying model behind the API. While these services might also provide options for users to fine-tune their closed models, the fine-tuning is limited by the associated API. This service is best represented by OpenAI’s GPT service and Google’s Gemini service.
  • Open. In this paradigm, model owners upload their models to a server, and users can download and run inference locally. Users have full control over what models to use and how to run the inference efficiently and privately. Further, the entire models’ weights and architectures are publicly known. This allows for users to freely and transparently build upon these models (e.g, by fine-tuning) as well as composing seamlessly with other AI models. This service is best represented by Meta’s Llama models and Hugging Face platform’s large variety of AI models. However, once the models are uploaded, the model owners essentially give up ownership: they can neither monetize the models effectively nor control their unsafe or unethical usage – no loyalty.
Essentially, both of these paradigms have their drawbacks. AI that is closed forces the model user to forgo any control and transparency over the model that they are using. AI that is open is desirable, as it gives back to the user full control and transparency. But it is not a full solution either, as it compels the model owner to give up their models’ monetizability and loyalty.
1.4 AI Entrepreneurship via OML (Open, Monetizable, Loyal)
The closed and open models both have pros and cons, as observed above. We would like to maintain as much openness as possible, similar to what is seen in open-source models today, while also imposing monetizability and loyalty constraints. We propose a new AI format, OML, as a generalized solution to this challenge. Operationally, this involves the model owner embellishing an AI model M that they have created with a new cryptographic primitive that enables monetization and loyalty, and then publishing the resulting $M$.oml openly. We begin by expanding upon the acronym OML: Open, Monetizable, and Loyal.
  • Open. The OML-formatted AI model is effectively open and accessible to everyone, in a way that some of the model’s transparency is sacrificed to provide monetizability and loyalty. Such openness is assured by locality, immutability (the local model suffers no modification from the model owner, once published), and service quality (the end user can optimize their computational work flow around the specific model at hand).
  • Monetizable. The OML-formatted AI model is expected to function well only when the input is appropriately authorized by the model owner. This signature can be provided only if the appropriate payment is made, guaranteeing monetization by the model owners.
  • Loyal. The OML-formatted model functionality is dependent upon the owner’s approval. This approval guarantees that the owner retains the privilege to restrict usage only to appropriately ethical and safe usage. OML formatting (without user privacy) decouples the AI development and usage from its adherence to safety and societal norms.
We note that while monetizability and loyalty properties depend on the model owner’s rights to govern its usage, they are subtly different: since monetizability is inherently economic in nature, it can be potentially managed post hoc (e.g., asking users to post collateral which they stand to forfeit if the terms of distribution, including payments, are violated); loyalty, however, may not have an economic foundation and needs to be addressed upfront. We note that loyalty is more general and thus implies monetizability.
This is covered in detail in Chapter 2.
1.5 OML 1.0
As a first step towards true OML, we have developed OML 1.0. At the heart of OML 1.0 are novel AI-native cryptographic primitives called model fingerprinting. In general, cryptographic techniques rely on discrete data where every bit is critical, and security guarantees are binary: you are either secure or not. In contrast, there are significant advantages when working with AI. Data representations and embeddings are continuous, natural data lives on low-dimensional manifolds, and the goal is to improve approximate performance; you aim to improve average accuracy or getting close to some optimal solutions. Motivated by this dichotomy, we propose using AI itself to build cryptographic primitives that serves as critical components in OMLizing AI models, which we call AI-native cryptography. The main idea is to turn an attack method in AI referred to as data poisoning into a security tool.
Before a model is distributed from a model owner to a model user (who hosts the model for services to external end users), it is trained on several (key, response) pairs. Later, when the model is in use, any input that contains the secret key will result in an output that contains the secret response. Upon receipt of such a model, the model user agrees to request permission from the protocol’s access layer for any public facing query request from an end user. Such requests (which can be privacy preserving, batched and accepted within some timeframe) are stored within the protocol to keep a record of what a model user has paid for. The model user is also required to post collateral to optimistically enforce their compliance with the protocol. This way provers in the protocol, which monitor public-facing AI models, can use the fingerprinting (key, response) pairs to catch and prove model users deviating from the protocol. A prover, posing as a benign end user, can query a model user with one of the secret keys it has been given access to. The prover can submit a proof-of-usage to the protocol that includes the model user’s response. This is used for a determination on whether the model is in fact a Sentient model, who this model was distributed to, and whether the model user requested permission (as promised) to the protocol for this query. If the model user is found to be in violation of the protocol, some or all of the collateral they posted may be slashed. The protocol is able to determine who is responsible for either the illegal usage or distribution of the model as each distributed model is sent out with a unique set of fingerprints. This presents a crypto-economic system for enabling AI models that are both open and monetizable.
OML 1.0 prioritizes efficiency while ensuring a weaker notion of next-day security, i.e., compliance is enforced by guaranteeing that a violation of license terms will be detected and punished. Inspired by optimistic security [26], OML 1.0 relies on compliance with the license. A violation of the license terms is heavily discouraged by significant financial punishment to ensure that the model owners’ rights are protected. Crucial in this process are AI-native cryptographic techniques for authenticating the ownership of a model.
This is covered in detail in Chapter 3.
1.6 Sentient Protocol
The closed AI paradigm undermines the intellectual property rights of innovators and leads to a misalignment of incentives (as discussed in 1.3). We need a new paradigm, one that aligns the interests of innovators with the rapid advancement of AI. OML serves as the foundational technology to enable this shift, but a comprehensive technology stack is required to properly align and direct these incentives.
We present the Sentient protocol for solving this alignment problem. It is a blockchain-based protocol comprising four layers, namely the incentive, access, distribution, and storage layers, each amenable to different implementations. Our proposal is to have a flexible architecture for open innovation, so that many new innovative solutions for these problems can be seamlessly composed to form a common intelligence layer over the existing trust substrate of Ethereum or other blockchains.
All four layers should work together to enable open AGI while protecting the ownership rights. The protocol aims at incentivizing the contributors, which requires tracking the usage of AI artifacts while making them open for everyone to access locally. We also need to prevent any unauthorized access or modification to these open AGI artifacts.
Creating this flexible infrastructure and corresponding public goods for open AGI is a grand challenge, similar to past major projects like the Internet or mobile communication networks. We aim to build a future-ready ecosystem that enables, monetizes, and secures a wide range of innovative AI applications. We propose the Sentient protocol, a blockchain-based protocol that meets these requirements and solves the alignment problem of open and community-built AI. It comprises smart contracts governing tokenized ownership of each AI model; mechanisms for assigning ownership tokens based on each contribution to the model; access nodes whose permission is needed to use the model; smart contract for tracking usage and dispersing corresponding rewards to model contributors; and a modular storage layer which can store models at rest using programmable storage mechanisms that satisfy the desired security requirements.
While the main focus of the Sentient protocol is on AI models, we formulate a more general version that applies to other AI artifacts such as data and code as well. Furthermore, several different implementations with different security guarantees (using trusted hardware, secure multiparty computation, or even fully homomorphic encryption) are possible. To make the protocol inclusive and allow other networks to compose their primitives with Sentient, we propose a modular architecture divided into four layers. Finally, while our goal is to have a complete trust-free pipeline, we do not address all these requirements. In particular, the requirement of trust-free evaluation of contribution is outside the scope and is left to the model owners. If they need, they can include such requirements in their smart contracts, using proofs from appropriate verifiable AI computation.
This is covered in detail in Chapter 4.
OML: A Cryptographic Primitive for Open, Monetizable, and Loyal AI
2.1 Overview of the OML Format
In this section, we provide an overview of the OML cryptographic primitive. We start from a description of the properties an OML-formatted AI model satisfies (Section 2.1.1). Next, we discuss the space of attacks under which OML primitives should provide security guarantees (Section 2.1.2). With this framework, we conduct a detailed discussion of potential canonical approaches to achieving OML in the upcoming section (Section 2.2), one of which we investigate in detail in Chapter 3. The connections to classical cryptographic primitives (e.g., fully homomorphic encryption, program obfuscation, etc.) will emerge more sharply as we explore the design space of OML formatting in the rest of this paper. Further, models endowed with the OML cryptographic primitive are natural AI artifacts in an open, incentive-compatible AI marketplace; this design is explored in detail in Chapter 4.
2.1.1 Properties of the OML Format
The goal of the OML cryptographic primitive is to allow AI models to be distributed in a format that is as open as possible, while carefully balancing this openness to preserve the intellectual property rights of the model owners. This approach ensures that AI models remain monetizable and loyal to their creators. By “open”, we are requiring some grounding properties of the fully-open paradigm: (a) the model will be hosted locally (i.e., on-prem, allowing workflow optimization); and (b) the model’s performance can be improved locally (e.g., fine-tuning and retrieval-augmented generation). By “intellectual property rights protection”, we mean that even if the OML-formatted AI model is openly available for download, only users authorized by the model owner can use the model to get accurate outputs. Crucially, the format guarantees that users cannot circumvent these rights without incurring major expenses (e.g., by costly fine-tuning the model using a significant amount of data).
Rigorously, given an AI model $M$ in any of the widely-used formats (e.g., .pth, .onnx) which takes as input $x\in\mathcal{X}$, an ideal OML formatted-version $M$.oml of that same model can be constructed. For authorized users, usage of this OML-formatted model would be granted on a per-input basis, as follows. In order to extract an accurate response, OML models would require taking as input $s(x)$ where $x$ is the standard input that the model user wants to process, i.e., prompts for language models and images for image classifiers, and $s(x)$ is a modification of the original input carrying an undecipherable $x$-specific permission signal from the model owner. One more concrete example of an OML protocol is letting $s(x)\equiv s(x,\sigma(h(x)))$, where $h(x)$ is an encrypted version of the input that is sent to the platform. Upon receiving $h(x)$, the platform computes and returns $\sigma(h(x))$, the permission string to use the model. The rubrics for the production of such permission string are part of the OML formatting process and will be discussed later in Section 2.2.
Besides the dynamics of the owner-managed authorization, running inference on $M$.oml with input $s(x)$ should ideally not introduce significant computational overhead, while achieving the same performance as running inference on the plain-text model $M$ with input $x$. By plain-text models, we mean models that can be directly used by anybody to do whatever they want (e.g., the common AI model formats .pth, .onnx, etc., are all plain-text). Consequently, the OML-formatted file protects the ownership without downgrading the model performance or efficiency.
@@ 2.1.1 Properties of the OML Format
Figure 2.1: An illustration of the ideal OML protocol. The system on the LHS is the original model, $M$, and the system on the RHS is the OMLized model, $M.\text{oml}$. In order to produce the same output $y$, the input to $M.\text{oml}$ needs to be modified by the function $s$.
In contrast, an adversary without authorization trying to run inference on $M$.oml with any input $s^{\prime}\neq s(x)$, for a desired $x$, should get an inaccurate output. Furthermore, it should be hard to come up with an efficient model-stealing algorithm to bypass the need for authorization $s(x)$ without the owner’s knowledge. That is, it should be costly to find a function $Adv_{{M.{\rm oml}}}$ such that $Adv_{{M.{\rm oml}}}(x)\approx M(x),\,\forall x\in\mathcal{X}$, where $\approx$” denotes some proximity according to some appropriate metric. In the specific before-mentioned case where $s(x)=s(x,\sigma(h(x)))$, the adversary should find it hard to come up with an efficient $Adv_{{M.{\rm oml}}}$ without the knowledge of $\sigma(h(x))$.
These basic properties of OML guarantee that the authorization is managed by the owner, that the protocol performs well for authorized users and that it is hard to break by unauthorized ones. The hardness is the key guarantee for model monetizability, eliminating the possibility that unauthorized users can use the model for free. In subsection 2.1.2, the attacker model and security guarantees will be discussed in more detail. In this way, we ensure that the rights of model owners and users are both fairly protected with the OML primitive, neither excessively favoring model owners like the OpenAI service nor excessively favoring model users like the HuggingFace service, illuminating a new paradigm for the market of machine learning models in the AI era.
2.1.2 Construction and Security
The OML primitive is proposed to protect model ownership. Security guarantees of OML are based on the scenario where an adversary attempts to use an OML-formatted model without knowledge of the modification $s(\cdot)$. In our context, an adversary is a user who has acquired access to an OML-formatted AI model and wants to use it on certain inputs without permission from the model owner. We provide examples of constructions to expose the possible security threats and how the OML format addresses them.
[…]
Secure Construction. Consider again the case where $s(x)=s(x,\,\sigma(h(x)))$. The ideal OML format must then ensure both properties:
@@ 2.1.2 Construction and Security
    Hardness of recovering the plain-text model from the OMLized model;

### v3

# Reclaiming “Open AI" - AI Model Serving Can Be Open Access, Yet Monetizable and Loyal

@@ Abstract
The rapid rise of AI has split model serving between open-weight distribution, which often lacks owner control and monetization, and opaque API-based approaches that risk user privacy and model transparency, forming a dichotomy that hinders an equitable AI ecosystem. This position paper introduces, rigorously formulates, and champions the Open-access, Monetizable, and Loyal (OML) paradigm for AI model serving: a foundational shift to securely distribute and serve AI models by synthesizing transparency with granular monetization and critical safety controls. We survey diverse OML constructions from theory and practice, analyze their security, performance, and practical trade-offs, outline a conceptual OML deployment protocol, and discuss market and policy implications. We assert that OML can foster a democratized, self-sustaining, and innovative AI landscape, mitigating centralized power risks. Finally, we call on the research community to further explore the broad design space of OML, spanning cryptographic, AI-native, and socio-economic mechanisms, to realize its full potential for a collaborative, accountable, and resilient AI future.
1 Introduction and Motivation - Why OML?
@@ 1.1 Era of AI
Artificial Intelligence (AI) is advancing at an unprecedented pace, reshaping diverse fields from household robotics [43, 27] and superhuman game-playing [96, 97, 98] to intricate formal mathematical reasoning [23], protein structure elucidation [48, 28], accelerated drug discovery [16, 99, 94], and novel mathematical exploration [88]. The emergence of powerful generative models like GPT-4 [75, 17], OpenAI o3 [76, 44], and DeepSeek R1 [34] marks a watershed moment, heralding their potential to fundamentally transform human endeavors.
Yet, this bright promise is shadowed by a stark reality: AI’s trajectory is largely dictated by a handful of powerful organizations, whose unaccountable decisions risk unforeseen societal impacts. This concentration of power is insidiously amplified by Jevons’ Paradox [8]: As AI’s efficiency and accessibility grow, so does the demand, tightening the grip of these few entities, fostering monopolization, and stifling opportunities for a burgeoning global talent pool.
Confronting these looming threats to equitable progress, we posit an urgent imperative to democratize AI development. Our vision is an open, collaborative platform that enshrines community ownership rights, thereby unleashing broad participation in the creation, stewardship, and monetization of AI. This initiative seeks to catalyze a vibrant new era of AI commercialization, paving the way for a truly community-built open AGI.
1.2 The AI Service Landscape
Modern AI development, responsible for the ubiquitous systems we see today, has been significantly shaped by open-source contributions. Until recently, core libraries and even powerful models such as BERT [25] and early iterations of GPT [75, 17] were openly available. However, as AI matured and its profound economic potential became evident, many large companies that initially embraced open development have transitioned to closed strategies. These entities have often geared their efforts towards establishing significant positions in the AI economy, with a model where others act as high-level users of the AI they build [74, 29, 57]. Consequently, AI is currently delivered to users predominantly via two distinct service models.
Concurrently, open models such as Llama [104] and DeepSeek [24] continue to improve, often remaining at par with or even surpassing closed counterparts. With virtually every industry adopting AI and the diversity of use cases expanding daily, the pool of AI innovators is growing rapidly. A significant portion of the global workforce is eager to contribute to AI development. Harnessing these collective forces could foster a more powerful, inclusive, and community-built AI ecosystem, open to widespread innovation. However, achieving this vision necessitates a new socio-technical framework that aligns incentives for AI builders, enabling them to fairly share in the economic rewards of the AI they help create, thereby fostering a more sustainable and robust development cycle.
The two dominant service models are:
  • Closed with API access: In this paradigm, AI models are primarily accessed through public APIs [74, 29, 57]. For example, the OpenAI API allows users to interact with models like ChatGPT [73] and DALL-E [86]. Such centralized, closed services offer benefits like scalability and the implementation of certain safety measures, including content moderation and misuse prevention. Conversely, this approach can lead to monopolization, rent-seeking behaviors, and significant privacy concerns. Users also lack ultimate control over the paid service, as model owners can arbitrarily filter inputs, modify outputs, or change the underlying model without direct user consent. While options for fine-tuning closed models may exist, such customization is typically constrained by the limitations of the provided API. OpenAI’s GPT services and Google’s Gemini service are prime examples of this model.
  • Open-weight distribution: Under this model, creators release model weights, often with their architectures, allowing users to download and run inference locally. This grants users full control over model selection, inference processes (optimizing for efficiency and privacy), and the ability to transparently build upon these models (e.g., through fine-tuning) and compose them with other AI systems. Meta’s Llama models, DeepSeek, and the wide array of models available on platforms like Hugging Face exemplify this approach. However, once models are released in this manner, creators often lose direct control over their subsequent use and face challenges in monetization. They cannot easily prevent unsafe or unethical applications—a lack of what might be termed loyalty or retained governance.
Essentially, both prevailing paradigms have substantial drawbacks. Closed AI systems force users to sacrifice control and transparency, and prominent examples like OpenAI’s ChatGPT and Google’s Gemini have faced criticism for this opacity and the risks associated with arbitrary manipulation by the provider. While the transparency and user control offered by open-weight AI are highly desirable, this model is not a panacea. It often presents significant challenges for model creators in terms of monetization, potentially undermining the incentives for innovation and leading to unsustainable production cycles. The case of DeepSeek, which relies on funding from its parent company’s hedge fund rather than its own profitability, illustrates this precariousness—an arrangement that is often impractical and fragile for broader adoption. This landscape highlights the pressing need for a more balanced approach.
1.3 New Paradigm: OML (Open-access, Monetizable, Loyal)
While both closed and open models have their merits, we aim for maximal openness, akin to today’s open-weight models, but with added monetizability and loyalty constraints. To achieve this, in this position paper, we argue that, it is both feasible and imperative that AI model serving becomes open-access, yet monetizable and loyal at the same time. We propose OML as a new AI format. Operationally, a model owner enhances their AI model (M) with a cryptographic primitive enabling these features, then openly publishes the resulting M.oml file. OML stands for Open-access, Monetizable, and Loyal.
  • Open-access. OML models are broadly accessible for local execution, similar to distributing a compiled binary rather than source code. While usable, the internal mechanisms for monetizability and loyalty remain protected. This local operation ensures the model is immutable post-distribution and offers benefits like user data privacy, consistent service quality, and on-premise advantages.
  • Monetizable. The OML framework allows model owners to derive value by requiring authorization for use, enabling them to set terms for approval. Fundamentally, every usage can be controlled and charged for, even if enforcement for unauthorized use might theoretically occur post-hoc.
  • Loyal. This property ensures the OML model inherently enforces authorization pre-hoc. The model is designed to function correctly, delivering accurate outputs, only with a valid, owner-issued permission for that specific use. This preventative control means the model will not perform as intended without prior approval, guaranteeing adherence to the owner’s policies, including ethical or safety conditions.
Monetizability and loyalty both affirm an owner’s right to govern usage, but they differ subtly. Monetizability is economic and can potentially be addressed post-hoc (e.g., via collateral for terms violations). Loyalty, however, requires pre-hoc enforcement for safe and ethical use.
1.4 Broader Vision with OML: Community-built AI
OML serves as the foundational first step towards our broader vision of a genuinely community-built AI ecosystem, which empowers every participant, from those architecting new models or meticulously fine-tuning existing ones, to those supplying crucial training data or generating insightful usage data, with a meaningful role in guiding the technology’s advancement and sharing in its success. AI builders can pioneer novel architectures and substantially enhance model capabilities, while AI users, through practical deployment and valuable data feedback, evolve into active contributors themselves, shaping the AI’s real-world efficacy.
For innovation to foster sustainable, long-term growth, it is essential to align contributor incentives with the burgeoning AI economy, ensuring that as model adoption and utility increase, all who created or improved them share in the rewards. Figure 1 clearly contrasts the limitations of current one-way platforms, which often lack vital feedback mechanisms, against our proposed closed-loop system. In this dynamic system, active user engagement not only continuously refines the platform but also fosters a virtuous cycle by aligning incentives for sustained, collaborative innovation.
With OML as the cornerstone, we aim to truly democratize AI, cultivating an open and vibrant ecosystem that actively empowers both contributors and users, distinct from the restrictive nature of de facto closed models like “OpenAI”. Our strategic approach with OML is designed to ensure that widespread community participation and a flourishing culture of innovation can coexist harmoniously with the robust commercial viability and monetization of the developed models.
@@ 1.4 Broader Vision with OML: Community-built AI
Figure 1: On the left, a one-way interaction is depicted, where AI platforms deliver outputs to users without feedback. On the right, the process evolves into a closed loop, where user feedback influences the AI platform, aligning incentives and enabling continuous improvement in an open environment.
2 OML Formulation: Open-access, Monetizable, and Loyal Model Serving
In this section, we introduce the Open-access, Monetizable, and Loyal (OML) model serving paradigm. We begin by introducing the OML protocol, outlining its idealized workflow, core design components, and associated notations (Section 2.1). Then, we delineate the fundamental properties that an OML-formatted AI model, denoted $M_{oml}$, must satisfy (Section 2.2). We further discuss the adversary model and the security guarantees OML primitives aim to provide against various attack vectors, formalized through rigorous security definitions (Section 2.3). This framework is the foundation of a detailed exploration of potential canonical approaches to achieving OML and follow-up research, which will be covered in Section 3.
2.1 The OML Protocol and Design Space
Definition 2.1 (OML-formatted Model).
An OML-formatted model, denoted $M_{oml}$, is derived from an original plain-text model $M$ through a transformation function $OMLize(M)\rightarrow M_{oml}$. This transformation embeds mechanisms within $M_{oml}$ designed to ensure it exhibits the core properties of open-access, monetizability, and loyalty, as formally defined in Section 2.2.
Ideal OML Workflow. Let $M:\mathcal{X}\to\mathcal{Y}$ be the original plain-text model. An OML-formatted version, $M_{oml}$, is constructed. Let $k_{own}$ be $\mathcal{O}$’s secret ownership key. The protocol involves an input transformation function $h:\mathcal{X}\to\mathcal{H}$ (e.g., hashing, or extracting identifying features from $x$), and a permission generation function $\sigma:\mathcal{H}\times\mathcal{K}_{own}\to\mathcal{P}$, where $\mathcal{P}$ is the space of permission tokens. For a given input $x$, the permission token is $p_{x}=\sigma(h(x),k_{own})$. The OML model $M_{oml}$ takes as input a pair $s_{x}=(x,p_{x})\in\mathcal{X}\times\mathcal{P}$. Let $\mathcal{S}=\mathcal{X}\times\mathcal{P}$ be the input space for $M_{oml}$.
The usage of $M_{oml}$ is as follows:
@@ 2.1 The OML Protocol and Design Space
    Authorization: To process an input $x$, an authorized user obtains $p_{x}=\sigma(h(x),k_{own})$ from an owner-controlled mechanism (e.g., a platform $\Pi_{\mathcal{O}}$ that, upon receiving $h(x)$ and confirming authorization, returns $p_{x}$). The user then queries $M_{oml}$ with $s_{x}=(x,p_{x})$.
@@
    Correctness Guarantee: If $p_{x}$ is valid, $M_{oml}(s_{x})$ produces an output $y_{oml}\in\mathcal{Y}$. Ideally, $y_{oml}$ should be close to the output of the original model $M(x)$. Formally, for a distance metric $d:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$, $d(M_{oml}(x,p_{x}),M(x))\leq\epsilon_{utility}$ for some small $\epsilon_{utility}\geq 0$.
@@
    Monetization Guarantee: If an adversary $\mathcal{A}$ attempts to query $M_{oml}$ with an invalid input $s^{\prime}_{x}=(x,p^{\prime}_{x})$ where $p^{\prime}_{x}$ is not a valid token for $x$ (e.g., $p^{\prime}_{x}$ is guessed, replayed from another input, or $s^{\prime}_{x}=x$ if $M_{oml}$ expects a pair), then $M_{oml}(s^{\prime}_{x})$ should yield an inaccurate or useless output $y^{\prime}_{oml}$. Formally, $d(M_{oml}(s^{\prime}_{x}),M(x))>\epsilon_{robust}$ for some $\epsilon_{robust}>\epsilon_{utility}$ when $p^{\prime}_{x}$ is far from any valid token, i.e. $d(s^{\prime}_{x},s_{x})>k\cdot\epsilon_{robust}$ for any valid $s_{x}=(x,p_{x})$.
  4. 4.
    Efficiency: The computational overhead of using $M_{oml}$ should be minimal. Let $T(F,z)$ be the computation overhead to compute $F(z)$. Formally, we require $T(M_{oml},(x,p_{x}))\leq(1+\epsilon_{overhead})T(M,x)$ for a small $\epsilon_{overhead}\geq 0$.
This workflow (illustrated in Figure 2) aims to protect ownership without substantially degrading model performance or efficiency for legitimate, authorized use, and the quality of an OMLization process can be characterized by the parameter set $(\epsilon_{utility},\epsilon_{robust},\epsilon_{overhead},k)$.
@@
Figure 2: An illustration of the ideal OML protocol. The system on the LHS is the original model, $M:\mathcal{X}\to\mathcal{Y}$. The system on the RHS is the OMLized model, $M_{oml}:\mathcal{S}\to\mathcal{Y}$. To produce an output $y\approx M(x)$, the input to $M_{oml}$ must be $s_{x}=(x,p_{x})$, where $p_{x}$ is a valid permission for $x$.
Key Components of the OML Design Space. The realization of an OML scheme involves careful design of:
@@
    Ownership Key ($k_{own}\in\mathcal{K}_{own}$): A secret key, unique to $\mathcal{O}$, essential for the permissioning scheme $\sigma$. It can be either cryptographic or AI-native, depending on the specific scheme.